OpenStudy (zab505):

If a polynomial is divided by (x - a) and the remainder equals zero, then (x - a) is a factor of the polynomial. A. True B. False

OpenStudy (anonymous):

True :)

OpenStudy (anonymous):

And to take that one step further, if you solve for a: x - a = 0 x - a + a = 0 + a x = a that (the opposite of a; in this case, it's positive a instead of negative) is a solution of the function.